OpenStudy (anonymous):

Generate the first 5 terms of this sequence: f(1) = 2 and f(2) = 3, f(n) = f(n - 1) + f(n - 2), for n > 2. A. 2, 3, 5, 7, 9 B. 2, 3, 4, 7, 11 C.2, 3, 5, 9, 11 D. 2, 3, 5, 8, 13

OpenStudy (cwrw238):

for n greater than 2:- 3rd term = f(3) = f(3-1) + f(3-2) = f(2) + f(1) = 3 + 2 = 5 continue with this process to find 4th 5th terms, 6th terms

OpenStudy (anonymous):

can you help me do the others please @cwrw238

OpenStudy (cwrw238):

* 4th snd 5th terms